# Look-ahead Search on Top of Policy Networks in Imperfect Information Games

### v2

@@ 3.2 Regularized Nash Dynamics
Using usual reinforcement learning techniques to find an optimal policy with self-play does not guarantee convergence to Nash equilibrium and can be exploited (Lanctot et al., 2017). However, regularizing the reward leads to establishing this guarantee (Perolat et al., 2022; Sokota et al., 2023).
@@
where $\pi^{R}$ is a regularization policy and $\eta$ is a hyperparameter for adjusting the regularization effect. Games with such a regularized reward have a single unique fixed point $\pi_{\text{fix}}$ for any regularization policy $\pi^{R}$ (Perolat et al., 2021). Follow the Regularized Leader (McMahan, 2011) may be used to converge to $\pi_{\text{fix}}$. Setting $\pi^{R}=\pi_{\text{fix}}$ will change the unique fixed point. If this is done repetitively, the fixed point will gradually converge to the Nash equilibrium.
In order to converge to a Nash equilibrium in large games with reward regularization, the authors in (Perolat et al., 2022) propose using two-player V-trace with Neural Replicator Dynamics (Hennes et al., 2020) as a loss function for policy head and usual L2 loss for value head. With this method, it is possible to achieve human-level performance in Stratego.
@@ 4 Multi-agent Search with Policy Transformations
In this section, we present an algorithm Multi-agent Search with Policy Transformations (SePoT) that trains an additional critic alongside a policy network with some policy-gradient algorithm for imperfect information games, like RNaD (Perolat et al., 2022). During gameplay, this critic is used to construct a depth-limited gadget game that is then solved with some search algorithm, in our case CFR+ (Tammelin et al., 2015).
@@ 4.1 Network Training
Performing a depth-limited search in imperfect information games requires a value function that, after the depth limit, assigns value to each infoset based on ranges (Kovařík et al., 2023). To use a value function based on multi-valued states, it is necessary for each history $h$ to learn multiple values that correspond to the expected values if the opponent plays some fixed policies against the player’s blueprint policy. Training such value function presents two difficulties to overcome. The first is the selection of the opponent policies to be used as a basis to compute the expected values. In large games, simply using some predefined policy is not viable due to the sheer size of full policies. The second is estimating the expected value of these policies.
To find different opponent policies, we decided to use a similar method to the ”bias approach” in (Brown et al., 2018). The main idea is to use a single policy, which in our case is the policy from the policy network, and then apply some mapping that maps the network policy to a different policy. We call this mapping a policy transformation.
@@
In poker, an example of a transformation is folding more frequently (Bowling et al., 2009; Bard et al., 2014). Similarly, in Goofspiel, certain transformations increase the probability of playing cards with lower value if the player is winning or decrease the probability of playing even-valued cards. These transformations do not have to be game-domain specific. However, the effectiveness of different transformations varies across different games. These transformations aim to cover the strategy space as much as possible. Constant transformations that transform any policy into each pure strategy cover this space fully (Brown et al., 2018).
In our experiments, we use transformations that move the actor policy $\pi$ in some direction with a step size $k$, where $k\in\mathbb{R}$ is a constant. These transformations are encoded within a neural network, which undergoes concurrent training with the actor network. At each training step, for each trajectory, we compute the direction in which the actor policy for that trajectory has shifted. Subsequently, we update only the transformation associated with the minimum Euclidean distance across the trajectory. Since each update should improve the actor policy against the current opponent’s policy, this direction should aim into the important regions of strategy space.
@@
$\begin{split}&\mathcal{V}v_{i}(h)=v_{i}(h)+\\ &\mathbb{E}_{\mu}\left[\sum^{l}_{t=0}\rho^{t}\Big{(}\prod_{0\leq q<t}c^{q}\Big{)}\Big{(}R_{i}^{t}+v_{i}(h^{t+1})-v_{i}(h^{t})\Big{)}|h^{0}=h,\mu\right]\end{split}$ (2)
@@
(Inspired by (Espeholt et al., 2018)): Let $\rho^{t}=\min{(\overline{\rho},\frac{\pi(s_{i}(h^{t}),a_{i}^{t})}{\mu(s_{i}(h^{t}),a_{i}^{t})})}$, $c^{t}=\min{(\overline{c},\frac{\pi(s_{i}(h^{t}),a_{i}^{t})}{\mu(s_{i}(h^{t}),a_{i}^{t})})}$, $\overline{\rho}\geq\overline{c}\geq 1$ and $\mu(s,a)>0\;\;\forall s,a$. Let us assume there exists $\kappa\in(0,1]$, such that $\mathbb{E}_{\mu}\rho^{l}\Big{(}\prod_{{0\leq q<l}}c^{q}\Big{)}\geq\kappa$, where $l$ is a length of the trajectory. Then the V-trace operator $\mathcal{V}$ has a unique fixed point $v^{\pi_{\overline{\rho}}}$, which is the expected value of following policy
@@
Using $\overline{\rho}\rightarrow\infty$ will result in convergence to the $v^{\pi}$, that is, the expected value of the policy $\pi$ in given history. For training the history network, we use standard L2 loss.
@@ 4.2 Constructing public belief state
As discussed in the Section 2, naively employing search in the game could lead to highly exploitable strategies. Using resolving gadget (Burch et al., 2014), which limits the opponent’s counterfactual value in root infosets, bounds the error after performing search in a subgame. Constructing a gadget game requires the opponent’s counterfactual values $v^{\pi}_{o}$ in each root infoset and public belief state $\beta=(s_{0},r)$. Most previous methods stored counterfactual values $v^{\pi}_{o}$, public state $s_{0}$ and ranges $r$ from previous searches (Moravčík et al., 2017; Schmid et al., 2023). SePoT could leverage the same idea. Still, we propose a novel approach of using the trained history network to compute counterfactual values, the trained policy network to retrieve ranges for searching player $i$, and a function that reconstructs the public state. This allows running the search from an arbitrary public state without the need to search in all parent public states.
Safe search in imperfect information games has to be performed from all possible histories based on public information. This requires all histories $\mathcal{H}(s_{0})$ from current public state $s_{0}$ to construct the gadget game $\mathcal{G}^{G}_{i}$. Let us assume a function that generates all histories in the public state $s_{0}$ from the publicly available information to all players. While the details of reconstructing the public state $s_{0}$ are not explored in detail here, for games like Goofspiel or Battleships, the public state can potentially be reconstructed using Constraint Satisfaction Problem (CSP) (Russell and Norvig, 2009; Seitz et al., 2021).
Gadget game $\mathcal{G}^{G}_{i}$ requires ranges for the player $i$ and the chance player. Chance player ranges $r_{c}$ can easily be stored during gameplay like older methods (Moravčík et al., 2017; Schmid et al., 2023; Brown and Sandholm, 2018). Suppose the player $i$ played only by using the policy from the neural network during current gameplay and decides to perform a search at some public state $s_{0}$. In that case, it has to recompute the ranges $r_{i}$ for each infoset $s_{i}\in\mathcal{S}_{i}(s_{0})$. This requires backtracking from all of these infosets to all previous decision points and acquiring the policy from the policy network for each of them. Because we are in a perfect recall setting, each infoset cannot have more parenting infosets than the maximal trajectory length $l^{\text{max}}$ to this infoset from the initial state $w^{0}$, so this backtracking involves at most $|\mathcal{S}_{i}(s_{0})|l^{\text{max}}$ queries to the policy network.
@@ 4.3 Value Function
(b) Single public state
@@
Multi-valued states (Brown et al., 2018) are a modification of the original game, where after the depth limit, the opponent has one more decision point, in which it chooses its strategy for the remainder of the game against some fixed blueprint policy. This modified game is a standard game and can be solved by any game solving algorithm unaware of value functions. Authors of (Moravčík et al., 2017) trained value function for Poker, which served as a generalization of the value function from perfect information games. This value function returns counterfactual values based on ranges at the depth limit. These were then specially adapted for Counterfactual Regret Minimization (CFR) (Zinkevich et al., 2007).
These two approaches to the depth-limited search may not seem connected, but the authors of (Kovařík et al., 2023) show this connection by explaining that the multi-valued states could be viewed as a value function, which is not trained for all possible strategies in the game, but only for some subset. However, their explanation does not explicitly clarify how this connection could be leveraged.
@@
In the same example, using multi-valued states as proposed in (Brown et al., 2018) gives the opponent, player 2, another choice to pick between two strategies in each infoset, as shown in Figure 1(b).
@@
With the constructed gadget game and a corresponding value function, any previously developed search algorithm for imperfect information games may conduct the subsequent search. During our experiments, we used CFR+ (Tammelin et al., 2015) exclusively. When using CFR+, a subclass of Growing tree-CFR (GT-CFR) introduced in (Schmid et al., 2023), the bounds already established by the GT-CFR authors in Theorem 3 also apply to our approach. However, the quality of this value function $\xi$, which is the distance between values returned by the optimal value function and the trained value function, is affected by additional errors, even if the value function is trained perfectly. The first of these errors is that the blueprint strategy of resolving player $i$ is not Nash equilibrium but just some approximation. Second is that the opponent’s strategies used for multi-valued states may not be best responses to any strategy.
The multi-valued states could be expanded by giving both players a choice to pick from multiple strategies for the rest of the game (Brown and Sandholm, 2019; Kovařík et al., 2023). This would increase the number of values that must be trained quadratically because each pair of strategies between players yields different values. SePoT is compatible with this approach, and the only required change is to compute a Nash equilibrium of the underlying normal-form game at the depth limit during search instead of a best response.
We have already shown how the history critic trained for multi-valued states can be used as a value function. However, we could go one step further and train the usual value function out of this history critic. The value function may be trained before the gameplay phase by randomly sampling possible ranges, computing the counterfactual values from multi-valued states, and then training the value function on these counterfactual values. Such trained value function can then be used in already developed algorithms (Moravčík et al., 2017; Schmid et al., 2023) without any changes.
@@ 5 Experiments
We have conducted two experiments to validate SePoT’s performance. Our experiments have used the framework OpenSpiel (Lanctot et al., 2019). The depth-limited search was conducted with CPU Intel Xeon Scalable Gold 6146, operating at a frequency of 3.2 GHz, while the network training additionally used a single GPU Tesla V100. The hyperparameter setting is described in Appendix D. Appendix B provides a detailed description of the game rules.
@@ 5.2 Exploitability in smaller games
In the second experiment, we aim to evaluate the effect of depth-limited search with multi-valued states on the quality of the policy compared to the policy network. In the experiments, we compare the exploitability of the policy network from RNaD and the SePoT. We use Leduc hold’em and Goofspiel with 5 cards and randomized order of point cards for evaluation. We trained ten networks with different random seeds while computing exploitability every 500 iterations up to iteration 75,000. The results in Figure 3 show the mean and 95% confidence interval of the exploitability throughout these runs. In Leduc hold’em, we separately solve part before dealing a public card. In Goofspiel, we use different depth limits for search, where DL=$n$ means that both players play $n$ actions before the value function is called from the root of the current subgame.
@@
(a) Leduc hold’em
(b) Goofspiel
@@ 6 Related work
Previous search algorithms require using safe search both in train and test time (Brown et al., 2020; Brown and Sandholm, 2018; Moravčík et al., 2017; Schmid et al., 2023). Using those algorithms in larger games requires some subgame reduction techniques like abstracting the game (Brown and Sandholm, 2018) or revealing part of the private information (Zhang and Sandholm, 2021). SePoT uses a policy-gradient training algorithm that trains in arbitrarily sized games. Furthermore, SePoT chooses whether the search should be used at each point independently during gameplay. This enables safe search in larger games than previously possible without using any subgame reduction techniques. However, SePoT is flexible enough to be used with these subgame reduction techniques to increase the number of subgames that the search may be applied in even further.
@@
Our method is heavily based on the (Brown et al., 2018), with the notable differences being a more principled way to train the blueprint policy with policy-gradient algorithms and choosing the policy transformations. Moreover, we show how to use multi-valued states as a value function practically. We present a way to leverage an already trained critic network to construct a gadget game anywhere in the game without the need to perform a search in previous parts.
Model-free deep reinforcement learning methods based on regret minimization (Steinberger et al., 2020; Gruslys et al., 2020; McAleer et al., 2023) use search in the gameplay and still can be scaled to very large games. The main limitation of these methods is the need to compute the average strategy, which either requires storing all the trajectories used in learning or approximating this average, which is prone to error. In games like Battleships, the training may consist of billions of trajectories to train properly. Storing all of these trajectories and computing averages is both memory and computationally demanding. SePoT uses a policy-gradient algorithm that approximates the optimal strategy in immediate strategies, so it does not compute the average strategy.
@@ 7 Conclusion
Policy-gradient algorithms proved to be powerful algorithms for solving huge problems in various settings. Recently, these algorithms were successfully applied to two-player zero-sum imperfect information games and outperformed human professionals in Stratego. However, these algorithms could not incorporate search to improve mistakes caused by network inaccuracies. We have presented an algorithm SePoT that can be used alongside any trajectory sampling policy-gradient algorithm in two-player zero-sum imperfect information games to train additional critic network without any search during training. We have shown that this network is sufficient to enable safe search during test time. The search we use differs from previous approaches since it is independent of the application of search in previous parts of the game, with the only limitation being the local size of the subgame. Thanks to this, SePoT can employ search even in larger games while still having a choice to follow policy from the policy network. Our experimental results demonstrate that SePoT is able to outperform RNaD in most of the tested scenarios.

### v3

@@ 3.2 Regularized Nash Dynamics
Using usual reinforcement learning techniques to find an optimal policy with self-play does not guarantee convergence to Nash equilibrium and can be exploited Lanctot et al. (2017). However, regularizing the reward leads to establishing this guarantee Perolat et al. (2022); Sokota et al. (2023).
@@
where $\pi^{R}$ is a regularization policy and $\eta$ is a hyperparameter for adjusting the regularization effect. Games with such a regularized reward have a single unique fixed point $\pi_{\text{fix}}$ for any regularization policy $\pi^{R}$ Perolat et al. (2021). Follow the Regularized Leader McMahan (2011) may be used to converge to $\pi_{\text{fix}}$. Setting $\pi^{R}=\pi_{\text{fix}}$ will change the unique fixed point. If this is done repetitively, the fixed point will gradually converge to the Nash equilibrium.
In order to converge to a Nash equilibrium in large games with reward regularization, the authors in Perolat et al. (2022) propose using two-player V-trace with Neural Replicator Dynamics Hennes et al. (2020) as a loss function for policy head and usual L2 loss for value head. With this method, it is possible to achieve human-level performance in Stratego.
@@ 4 Multi-Agent Search With Policy Transformations
In this section, we present an algorithm Multi-Agent Search With Policy Transformations (SePoT) that trains an additional critic alongside a policy network with some policy-gradient algorithm for imperfect information games, like RNaD Perolat et al. (2022). During gameplay, this critic is used to construct a depth-limited gadget game that is then solved with some search algorithm, in our case CFR+ Tammelin et al. (2015).
@@ 4.1 Network Training
Performing a depth-limited search in imperfect information games requires a value function that, after the depth limit, assigns value to each infoset based on ranges Kovařík et al. (2023). To use a value function based on multi-valued states, it is necessary for each history $h$ to learn multiple values that correspond to the expected values if the opponent plays some fixed policies against the player’s blueprint policy. Training such value function presents two difficulties to overcome. The first is the selection of the opponent policies to be used as a basis to compute the expected values. In large games, simply using some predefined policy is not viable due to the sheer size of full policies. The second is estimating the expected value of these policies.
To find different opponent policies, we decided to use a similar method to the ”bias approach” in Brown et al. (2018). The main idea is to use a single policy, which in our case is the policy from the policy network, and then apply some mapping that maps the network policy to a different policy. We call this mapping a policy transformation.
@@
In poker, an example of a transformation is folding more frequently Bowling et al. (2009); Bard et al. (2014). Similarly, in Goofspiel, certain transformations increase the probability of playing cards with lower value if the player is winning or decrease the probability of playing even-valued cards. These transformations do not have to be game-domain specific. However, the effectiveness of different transformations varies across different games. These transformations aim to cover the strategy space as much as possible. Constant transformations that transform any policy into each pure strategy cover this space fully Brown et al. (2018).
In our experiments, we use transformations that move the actor policy $\pi$ in some direction with a fixed step size $k\in\mathbb{R}$. These transformations are encoded within a neural network, which undergoes concurrent training with the actor network. At each training step, for each trajectory, we compute the direction in which the actor policy for that trajectory has shifted. Subsequently, we update only the transformation associated with the minimum Euclidean distance across the trajectory. Since each update should improve the actor policy against the current opponent’s policy, this direction should aim into the important regions of strategy space, but the quality of these transformations is not guaranteed.
@@
$\begin{split}&\mathcal{V}v_{i}(h)=v_{i}(h)+\\ &\mathbb{E}_{\mu}\left[\sum^{l}_{t=0}\rho^{t}\Big{(}\prod_{\mathclap{0\leq q<t}}c^{q}\Big{)}\Big{(}R_{i}^{t}+v_{i}(h^{t+1})-v_{i}(h^{t})\Big{)}|h^{0}=h,\mu\right]\end{split}$ (2)
@@
(Inspired by Espeholt et al. (2018)): Let $\rho^{t}=\min{(\overline{\rho},\frac{\pi(s_{i}(h^{t}),a_{i}^{t})}{\mu(s_{i}(h^{t}),a_{i}^{t})})}$, $c^{t}=\min{(\overline{c},\frac{\pi(s_{i}(h^{t}),a_{i}^{t})}{\mu(s_{i}(h^{t}),a_{i}^{t})})}$, $\overline{\rho}\geq\overline{c}\geq 1$ and $\mu(s,a)>0\;\;\forall s,a$. Let us assume there exists $\kappa\in(0,1]$, such that $\mathbb{E}_{\mu}\rho^{l}\Big{(}\prod_{{0\leq q<l}}c^{q}\Big{)}\geq\kappa$, where $l$ is a length of the trajectory. Then the V-trace operator $\mathcal{V}$ has a unique fixed point $v^{\pi_{\overline{\rho}}}$, which is the expected value of following policy
@@
Using $\overline{\rho}\rightarrow\infty$ will result in convergence to the $v^{\pi}$, that is, the expected value of the policy $\pi$ in given history. The proof of Theorem 1 is in Appendix A. For training the history network, we use standard L2 loss.
@@ 4.2 Constructing Public Belief State
As discussed in the Section 2, naively employing search in the game could lead to highly exploitable strategies. Using resolving gadget Burch et al. (2014), which limits the opponent’s counterfactual value in root infosets, bounds the error after performing search in a subgame. Constructing a gadget game requires the opponent’s counterfactual values $v^{\pi}_{o}$ in each root infoset and public belief state $\beta=(s_{0},r)$. Most previous methods stored counterfactual values $v^{\pi}_{o}$, public state $s_{0}$ and ranges $r$ from previous searches Moravčík et al. (2017); Schmid et al. (2023). SePoT could leverage the same idea. Still, we propose a novel approach of using the trained history network to compute counterfactual values, the trained policy network to retrieve ranges for searching player $i$, and a function that reconstructs the public state. This allows running the search from an arbitrary public state without the need to search in all parent public states.
Safe search in imperfect information games has to be performed from all possible histories based on public information. This requires all histories $\mathcal{H}(s_{0})$ from current public state $s_{0}$ to construct the gadget game $\mathcal{G}^{G}_{i}$. Let us assume a function that generates all histories in the public state $s_{0}$ from the publicly available information to all players. While the details of reconstructing the public state $s_{0}$ are not explored in detail here, for games like Goofspiel or Battleships, the public state can be reconstructed by Constraint Satisfaction Problem (CSP) Russell and Norvig (2009); Seitz et al. (2021).
Gadget game $\mathcal{G}^{G}_{i}$ requires ranges for the player $i$ and the chance player. Chance player ranges $r_{c}$ can easily be stored during gameplay like older methods Moravčík et al. (2017); Schmid et al. (2023); Brown and Sandholm (2018). Suppose the player $i$ played only by using the policy from the neural network during current gameplay and decides to perform a search at some public state $s_{0}$. In that case, it has to recompute the ranges $r_{i}$ for each infoset $s_{i}\in\mathcal{S}_{i}(s_{0})$. This requires backtracking from all of these infosets to all previous decision points and acquiring the policy from the policy network for each of them. Because we are in a perfect recall setting, each infoset cannot have more parenting infosets than the maximal trajectory length $l^{\text{max}}$ to this infoset from the initial state $w^{0}$, so this backtracking involves at most $|\mathcal{S}_{i}(s_{0})|l^{\text{max}}$ queries to the policy network.
@@ 4.3 Value Function
((b)) Single public state
@@
Multi-valued states Brown et al. (2018) are a modification of the original game, where after the depth limit, the opponent has one more decision point, in which it chooses its strategy for the remainder of the game against some fixed blueprint policy. This modified game is a standard game and can be solved by any game solving algorithm unaware of value functions. Authors of Moravčík et al. (2017) trained value function for Poker, which served as a generalization of the value function from perfect information games. This value function returns counterfactual values based on ranges at the depth limit. These were then specially adapted for Counterfactual Regret Minimization (CFR) Zinkevich et al. (2007).
These two approaches to the depth-limited search may not seem connected, but the authors of Kovařík et al. (2023) show this connection by explaining that the multi-valued states could be viewed as a value function, which is not trained for all possible strategies in the game, but only for some subset. However, their explanation does not explicitly clarify how this connection could be leveraged.
@@
In the same example, using multi-valued states as proposed in Brown et al. (2018) gives the opponent, player 2, another choice to pick between two strategies in each infoset, as shown in Figure 2(b).
@@
With the constructed gadget game and a corresponding value function, any previously developed search algorithm for imperfect information games may conduct the subsequent search. During our experiments, we used CFR+ Tammelin et al. (2015) exclusively. When using CFR+, a subclass of Growing tree-CFR (GT-CFR) introduced in Schmid et al. (2023), the bounds already established by the GT-CFR authors in Theorem 3 also apply to our approach. However, the quality of this value function $\xi$, which is the distance between values returned by the optimal value function and the trained value function, is affected by additional errors, even if the value function is trained perfectly. The first of these errors is that the blueprint strategy of resolving player $i$ is not Nash equilibrium but just some approximation. Second is that the opponent’s strategies used for multi-valued states may not be best responses to any strategy.
The multi-valued states could be expanded by giving both players a choice to pick from multiple strategies for the rest of the game Brown and Sandholm (2019); Kovařík et al. (2023). This would increase the number of values that must be trained quadratically because each pair of strategies between players yields different values. SePoT is compatible with this approach, and the only required change is to compute a Nash equilibrium of the underlying normal-form game at the depth limit during search instead of a best response.
We have already shown how the history critic trained for multi-valued states can be used as a value function. However, we could go one step further and train the usual value function out of this history critic. The value function may be trained before the gameplay phase by randomly sampling possible ranges, computing the counterfactual values from multi-valued states, and then training the value function on these counterfactual values. Such trained value function can then be used in already developed algorithms Moravčík et al. (2017); Schmid et al. (2023) without any changes.
@@ 5 Experiments
We have conducted two experiments to validate SePoT’s performance. Our experiments have used the framework OpenSpiel Lanctot et al. (2019). The depth-limited search was conducted with CPU Intel Xeon Scalable Gold 6146, operating at a frequency of 3.2 GHz, while the network training additionally used a single GPU Tesla V100. The hyperparameter setting is described in Appendix D. Appendix B provides a detailed description of the game rules.
@@ 5.2 Exploitability in Smaller Games
In the second experiment, we aim to evaluate the effect of depth-limited search with multi-valued states on the quality of the policy compared to the policy network. In the experiments, we compare the exploitability of SePoT and the policy network from RNaD. We use Leduc hold’em and Goofspiel with 5 cards and randomized order of point cards for evaluation. We trained ten networks with different random seeds while computing exploitability every 500 iterations up to iteration 75,000. The results in Figure 3 show the mean and 95% confidence interval of the exploitability throughout these runs. In Leduc hold’em, we separately solve part before dealing a public card. In Goofspiel, we use different depth limits for search, where DL=$n$ means that both players play $n$ actions before the value function is called.
@@
((a)) Leduc hold’em
((b)) Goofspiel
@@ 6 Related Work
Previous search algorithms require using safe search both in train and test time Brown et al. (2020); Brown and Sandholm (2018); Moravčík et al. (2017); Schmid et al. (2023). Using those algorithms in larger games requires some subgame reduction techniques like abstracting the game Brown and Sandholm (2018) or revealing part of the private information Zhang and Sandholm (2021). SePoT uses a policy-gradient training algorithm that trains in arbitrarily sized games. Furthermore, SePoT chooses whether the search should be used at each point independently during gameplay. This enables safe search in larger games than previously possible without using any subgame reduction techniques. However, SePoT is flexible enough to be used with these subgame reduction techniques to increase the number of subgames that the search may be applied in even further.
@@
Our method is heavily based on the Brown et al. (2018), with the notable differences being a more principled way to train the blueprint policy with policy-gradient algorithms and choosing the policy transformations. Moreover, we show how to use multi-valued states as a value function practically. We present a way to leverage an already trained critic network to construct a gadget game anywhere in the game without the need to perform a search in previous parts.
Model-free deep reinforcement learning methods based on regret minimization Steinberger et al. (2020); Gruslys et al. (2020); McAleer et al. (2023) use search in the gameplay and still can be scaled to very large games. The main limitation of these methods is the need to compute the average strategy, which either requires storing all the trajectories used in learning or approximating this average, which is prone to error. In games like Battleships, the training may consist of billions of trajectories to train properly. Storing all of these trajectories and computing averages is both memory and computationally demanding. SePoT uses a policy-gradient algorithm that approximates the optimal strategy in immediate strategies, so it does not compute the average strategy.
@@ 7 Conclusion
Policy-gradient algorithms proved to be powerful for solving huge problems in various settings. Recently, these algorithms were successfully applied to two-player zero-sum imperfect information games and outperformed human professionals in Stratego. However, these algorithms could not incorporate search to improve mistakes caused by network inaccuracies. We have presented an algorithm SePoT that can be used alongside any trajectory sampling policy-gradient algorithm in two-player zero-sum imperfect information games to train additional critic network without any search during training. We have shown that this network is sufficient to enable safe search during test time. The search we use differs from previous approaches since it is independent of the application of search in previous parts of the game, with the only limitation being the local size of the subgame and reconstruction of the public states. Thanks to this, SePoT can employ search even in larger games while still having a choice to follow policy from the policy network. Our experimental results demonstrate that SePoT is able to outperform RNaD in most of the tested scenarios.